# Video Instance Shadow Detection

### v2

@@ Abstract
1 Introduction
Instance shadow detection, as exemplified by works [46, 44, 45], is a crucial task involving the identification of shadows alongside their associated objects in images. This task predicts shadow instances, object instances, and the relationships between them. Its applications span various domains, including photo editing and light direction estimation.
@@ 1 Introduction
To tackle these challenges, we introduce ViShadow, a semi-supervised framework for video instance shadow detection that leverages both labeled image datasets and unlabeled video sequences. Initially, we use labeled image data [46] to train ViShadow, enabling it to identify shadow instances, object instances, and their spatial relations. Subsequently, we employ center contrastive learning [14] to pair shadows/objects across different images. In the next phase, we leverage unlabeled videos, introducing an associated cycle consistency loss to enhance tracking ability by exploiting spatial relations and temporal information between shadow and object instances across frames. Lastly, we tackle the issue of temporary disappearance in videos, especially in occluded scenarios, through a retrieval mechanism, which enhances the tracking of shadow or object instances when portions of them vanish in specific frames. It is achieved by matching the learned tracking embedding for each shadow or object instance across various video frames, thereby improving the overall continuity of shadow-object associations throughout the video.
Additionally, we formulate the SOBA-VID dataset, comprising unlabeled videos for training and labeled videos for testing. Accompanying this, the SOAP-VID metric, an extension of the SOAP metric [46], facilitates the quantitative evaluation of VISD solutions. This metric, incorporating Intersection-over-Union (IoU) computation in a spatiotemporal context, allows for the assessment of the model’s performance in detecting and tracking shadow-object associations in video sequences.
The tracking masks derived from ViShadow, used to predict shadow-object associations, offer a valuable tool for various video editing tasks. For instance, as depicted in Fig. LABEL:fig:teaser, our framework enables the seamless extraction and insertion of a shadow-object pair from one video to another (or the same video). Additionally, VISD supports video inpainting [24, 29] and shadow editing, allowing for the obfuscation of entire objects by simultaneously addressing the object and its associated shadow, as illustrated in Figure 4. Moreover, leveraging the capabilities of current large vision-language models [1, 31], we achieve text-instructed shadow-object manipulation in videos. This involves interactive editing where user requirements are inputted through text commands, showcasing the potential of our framework for user-guided video editing.
@@
  • We integrate ViShadow into a platform capable of both mask-guided and text-guided shadow-object editing in videos. This showcases the adaptability of utilizing the results of video instance shadow detection across diverse tasks, including video inpainting, instance cloning, shadow editing, and text-instructed shadow-object manipulation.
2 Related Work
@@
2.1 Shadow Detection
Image shadow detection estimates a binary shadow region mask using methods ranging from physical techniques [38, 10, 34, 40], user input [2, 52], to handcrafted features [9, 22, 55, 13, 18]. Recent advancements in deep learning [17, 56, 53, 15, 6, 16, 50] have demonstrated remarkable performance. Extending this to video, Chen et al. [5] introduced ViSha and TVSD-Net, focusing on intra-video and inter-video feature learning. SC-Cor [8] improves temporal coherence, Lu et al. [33] addresses generalization errors, and Liu et al. [28] handles large shadow deformations. In contrast, our video instance shadow detection edits or removes individual shadows with associated objects at the video level, providing a unique perspective.
2.2 Instance Shadow Detection
Instance shadow detection identifies shadows and their associations with objects. LISA [46] pioneers this with a two-stage end-to-end framework for predicting bounding boxes, masks, and associations in images, featuring a strategy for pairing instances. SSIS [44, 45] adopts a single-stage fully convolutional network with bidirectional relation learning for understanding shadow-object relationships. The extension to video instance shadow detection encompasses both instance detection and paired tracking in videos. However, this extension introduces challenges not seen in image-based instance shadow detection, involving labor-intensive video annotation on a per-frame basis and tracking partial shadow-object pairs due the occlusion or out-of-view issues in videos, adding complexity to the task.
2.3 Video Instance Segmentation
@@
3 SEMI-Supervised Video Instance Shadow Detection Framework
@@
3.1 Overall Network Architecture
@@
In the following subsections, we elaborate on the methodology for acquiring a tracking embedding for instance-paired tracking (Section 3.2). Additionally, we introduce a bidirectional retrieving mechanism aimed at pairing up shadows or objects that have temporarily disappeared (Section 3.3).
@@
3.2 Instance Paired Tracking
3.2.1 Learning from Labeled Images
@@
3.2.2 Learning from Unlabeled Videos
In contrast to labeled image data providing instance labels, we utilize unlabeled videos for self-supervised learning in our ViShadow framework, as illustrated in Fig. 2 Stage II. For two consecutive frames in a video, we compute tracking embeddings of detected instances. Non-maximum suppression (NMS) is then applied, aiming to obtain distinct paired shadow-object tracking embeddings for each instance. Subsequently, a transition matrix $A$ is generated by computing the embedding similarity of instance pairs at different time frames. Each element $A_{t}^{t+1}(i,j)$ represents the transition probability from the $i^{th}$ instance at time $t$ to the $j^{th}$ instance at time $t+1$.
@@ 3.2.2 Learning from Unlabeled Videos
3.2.3 Tracking Paired Instances Across Frames.
@@
$\begin{split}Score(f_{i},f_{j})=&[\frac{exp(cos(f_{i},f_{j}))}{{\textstyle\sum_{k=1}^{M}}exp(cos(f_{k},f_{j}))}\\ &+\frac{exp(cos(f_{i},f_{j}))}{{\textstyle\sum_{k=1}^{N}}exp(cos(f_{i},f_{k}))}]/2\ ,\end{split}$ (5)
@@
3.3 Bidirectional Retrieving Mechanism for Unpaired Shadow/Object Instances
Table 1: Evaluation on the ViShadow framework design.
@@ 3.3 Bidirectional Retrieving Mechanism for Unpaired Shadow/Object Instances
4 Dataset and Evaluation Metrics
@@ 4.1 Dataset
SOBA-VID comprises 292 videos totaling 7045 frames, sourced from (i) existing datasets, including DAVIS [4], YouTube-VIS [51], and OVIS [35], (ii) self-collected videos, and (iii) Internet searches using the keyword “shadow plus common instances.” We randomly partition the videos into a training set (232 videos, 5863 frames) and a testing set (60 videos, 1182 frames). For the test set videos, we meticulously annotate frame-by-frame masks of each shadow/object instance using Apple Pencil, forming a total of 134 shadow-object pairs (refer to Fig. 5 for examples).
4.2 Evaluation Metrics
@@
Table 2: Comparison with two baseline methods for video instance shadow detection, showcasing our method’s substantial superiority. Improvements over “SSIS + Mask2Former” on SOAP-VID, Association AP, and Instance AP are 24.5%, 20.4%, and 20.6%, respectively.
@@
SSIS + IoU Tracker[3] 21.5 31.7 25.7
SSIS + Mask2Former[7] 31.8 51.1 42.2
@@
5 Experimental Results
5.1 Implementation Details
@@
For inference, ViShadow processes input videos online. Its tracking branch predicts tracking embeddings for both paired and unpaired shadow/object instances on a frame-by-frame basis, employing the matching strategy outlined in Section 3.2. Notably, ViShadow demonstrates its efficiency by completing a 21-frame video in 18 seconds.
5.2 Comparison with Baseline Methods
We compare our method with two baselines: (i) utilizing image-level detector SSIS [44] to detect the shadow instances, object instances, and shadow-object associations on each frame, and employing the IoU Tracker [3] to match instances across frames based on IoU values between detected bounding boxes; and (ii) employing the tracking algorithm Mask2former [7] to detect and track object instances across all video frames, using SSIS for shadow/object instance and shadow-object association detection on each frame, and merging the results based on mask IoU of each object instance.
Table 2 presents the comparison results, showing our method significantly outperforming both baselines across all evaluation metrics on the SOBA-VID test set. The IoU Tracker heavily relies on the quality of detection results and frame rate, while Mask2Former is constrained to recognizing objects of the same category as those in the training data [51]. In contrast, our approach benefits from end-to-end training and knowledge acquired from unlabeled videos. Visual comparisons between ViShadow and “SSIS + Mask2Former” are provided in Fig. 1.
@@ 5.2 Comparison with Baseline Methods
Table 3: Evaluation on the instance paired tracking design.
@@
5.3 Ablation Study
5.3.1 Component Analysis
We conduct an ablation study, incrementally introducing major components—tracking branch with contrastive learning in Eq.(3), temporal correspondence in Eq.(4), and bidirectional retrieving mechanism—starting from the SSIS baseline, which employs instance tracking through the mask controller’s output. The results in Table 1 demonstrate the contributions of each component, with contrastive learning and temporal correspondence enhancing paired tracking, and the retrieving mechanism addressing missed unpaired instances.
5.3.2 Instance Paired Tracking Analysis
We assess the effectiveness of instance paired tracking. Table 3 indicates that using tracking embeddings for individual shadow/object instances achieves paired tracking. And the best performance across all evaluation metrics is achieved by concatenating both shadow and object embeddings.
5.4 Comparison with OmnimatteRF
@@
6 Applications
6.1 Mask-Guided Shadow-Object Manipulation
@@ 6.1.3 Shadow Editing
6.2 Text-Instructed Shadow-Object Manipulation
@@
7 Conclusion and Limitations
@@
This paper introduces the video instance shadow detection task, aiming to detect, segment, associate, and track paired shadow-object instances in videos. We present the ViShadow framework, employing supervised learning on labeled images and self-supervised learning on unlabeled videos to learn paired tracking of shadow and object instances. A bidirectional retrieving mechanism is designed for robust paired tracking, even in cases where objects or shadows are temporarily missed. We contribute the SOBA-VID dataset for training and testing. Experimental results demonstrate the effectiveness of our approach across various metrics. Practical applications of video instance shadow detection in video-level editing scenarios are illustrated. However, this approach has certain limitations. First, distinguishing overlapped shadow instances from different objects is challenging because of their merged and indistinct boundaries. Second, accurately segmenting soft shadows, characterized by gradual edges and subtle intensity transitions, poses a significant challenge. Refer to Figure 10 for visual examples illustrating the challenges associated with overlapped and soft shadows. These constraints provide opportunities for future research and improvement.

### v3

@@ Abstract
Index Terms:
Instance shadow detection, shadow-object pairing, video analysis, and shadow detection.
Figure 1: Based on the video clip presented in (a), our ViShadow framework demonstrates a robust ability to detect, segment, associate, and track the dancer along with her shadow (b). This capability facilitates a range of applications, such as duplicating the dancer with her shadow to create captivating visual effects (c).
I Introduction
Instance shadow detection [1] is a crucial task involving the identification of shadows alongside their associated objects in images. This task predicts shadow instances, object instances, and the relationships between them by taking the single image as input. Its applications span various domains, including photo editing and light direction estimation.
@@ I Introduction
To tackle these challenges, we introduce ViShadow, a semi-supervised framework for video instance shadow detection that leverages both labeled image datasets and unlabeled video sequences. Initially, we use labeled image data [1] to train ViShadow, enabling it to identify shadow instances, object instances, and their spatial relations. Subsequently, we employ center contrastive learning [4] to pair shadows/objects across different images. In the next phase, we leverage unlabeled videos, introducing an associated cycle consistency loss to enhance tracking ability by exploiting spatial relations and temporal information between shadow and object instances across frames. Lastly, we tackle the issue of temporary disappearance in videos, especially in occluded scenarios, through a retrieval mechanism, which enhances the tracking of shadow or object instances when portions of them vanish in specific frames. It is achieved by matching the learned tracking embedding for each shadow or object instance across various video frames, thereby improving the overall continuity of shadow-object associations throughout the video. Following [1], we only detect the cast shadows produced by the corresponding objects and ignore the self-shadows on the objects themselves. Also, our method is designed specifically for scenarios involving a single light source, such as those under the sun and sky, and may not be directly applicable to environments with multiple light sources.
In contrast to our previous works [1, 2, 3], this paper not only detects shadows and their associated objects in video frames but also tracks each shadow, object, and their associations continuously throughout the entire video sequence, even when paired shadow or object instances temporarily disappear. Additionally, we formulate the SOBA-VID dataset, comprising unlabeled videos for training and labeled videos for testing. Accompanying this, the SOAP-VID metric, an extension of the SOAP metric [1], facilitates the quantitative evaluation of VISD solutions. This metric, incorporating Intersection-over-Union (IoU) computation in a spatiotemporal context, allows for the assessment of the model’s performance in detecting and tracking shadow-object associations in video sequences.
The tracking masks derived from ViShadow, used to predict shadow-object associations, offer a valuable tool for various video editing tasks. For instance, as depicted in Fig. 1, our framework enables the seamless extraction and insertion of a shadow-object pair from one video to another (or the same video). Additionally, VISD supports video inpainting [5, 6] and shadow editing, allowing for the obfuscation of entire objects by simultaneously addressing the object and its associated shadow, as illustrated in Figure 5. Moreover, leveraging the capabilities of current large vision-language models [7, 8], we achieve text-instructed shadow-object manipulation in videos. This involves interactive editing where user requirements are inputted through text commands, showcasing the potential of our framework for user-guided video editing. Our dataset, code, trained model, and evaluation metrics are publicly available at https://github.com/HarryHsing/Video-Instance-Shadow-Detection.
@@
  • We integrate ViShadow into a platform capable of mask-guided and text-guided shadow-object editing in videos. It showcases the adaptability of using the results of video instance shadow detection across diverse tasks, including video inpainting, instance cloning, shadow editing, and text-instructed shadow-object manipulation.
II Related Work
@@
II-A Shadow Detection
Image shadow detection estimates a binary shadow region mask using methods ranging from physical techniques [9, 10, 11, 12], user input [13, 14], to handcrafted features [15, 16, 17, 18, 19]. Recent advancements in deep learning [20, 21, 22, 23, 24, 25, 26, 27] have demonstrated remarkable performance. Extending this to video, Chen et al. [28] introduced ViSha and TVSD-Net, focusing on intra-video and inter-video feature learning. SC-Cor [29] improves temporal coherence, Lu et al. [30] addresses generalization errors, and Liu et al. [31] handles large shadow deformations. Please see [32] for the comprehensive survey on image and video shadow detection with deep learning. In contrast, our video instance shadow detection edits or removes individual shadows with associated objects at the video level, providing a unique perspective.
II-B Instance Shadow Detection
Instance shadow detection identifies shadows and their associations with objects. LISA [1] pioneers this with a two-stage end-to-end framework for predicting bounding boxes, masks, and associations in images, featuring a strategy for pairing instances. SSIS [2, 3] adopts a single-stage fully convolutional network with bidirectional relation learning for understanding shadow-object relationships. The extension to video instance shadow detection encompasses both instance detection and paired tracking in videos. However, this extension introduces challenges not seen in image-based instance shadow detection, involving labor-intensive video annotation on a per-frame basis and tracking partial shadow-object pairs due the occlusion or out-of-view issues in videos, adding complexity to the task. Please see [32] for the comprehensive survey and comparisons.
II-C Video Instance Segmentation
@@
III SEMI-Supervised Video Instance Shadow Detection Framework
@@
III-A Overall Network Architecture
@@
In the following subsections, we elaborate on the methodology for acquiring a tracking embedding for instance-paired tracking (Section III-B). Additionally, we introduce a bidirectional retrieving mechanism aimed at pairing up shadows or objects that have temporarily disappeared (Section III-C).
@@
III-B Instance Paired Tracking
III-B1 Learning from Labeled Images
@@
III-B2 Learning from Unlabeled Videos
In contrast to labeled image data providing instance labels, we utilize unlabeled videos for self-supervised learning in our ViShadow framework, as illustrated in Fig. 3 Stage II. Note that the center contrastive learning is not employed in Stage II due to the difficulty of accurately forming positive and negative pairs across frames in the absence of ground truth labels.
For two consecutive frames in a video, we compute tracking embeddings of detected instances. Non-maximum suppression (NMS) is then applied, aiming to obtain distinct paired shadow-object tracking embeddings for each instance. Subsequently, a transition matrix $A$ is generated by computing the embedding similarity of instance pairs at different time frames. Each element $A_{t}^{t+1}(i,j)$ represents the transition probability from the $i^{th}$ instance at time $t$ to the $j^{th}$ instance at time $t+1$.
@@ III-B2 Learning from Unlabeled Videos
III-B3 Tracking Paired Instances Across Frames.
@@
$\begin{split}Score(f_{i},f_{j})=&\frac{0.5\exp(\cos(f_{i},f_{j}))}{\sum_{k=1}^{M}\exp(\cos(f_{k},f_{j}))}\\ &+\frac{0.5\exp(\cos(f_{i},f_{j}))}{\sum_{k=1}^{N}\exp(\cos(f_{i},f_{k}))}\ ,\end{split}$ (5)
@@
TABLE I: Evaluation on the ViShadow framework design.
@@
III-C Bidirectional Retrieving Mechanism for Unpaired Shadow/Object Instances
During the training phase, our framework uses an association cycle consistency loss to ensure the consistency of tracking embeddings across frames. This loss helps maintain the identity of shadow-object pairs even when they temporarily disappear. In the testing phase, the bidirectional retrieval mechanism becomes crucial. If an instance (shadow or object) disappears in a frame, the mechanism stores its last known embedding. When the instance reappears, the retrieval mechanism uses the stored embeddings to re-associate the shadow and object, maintaining the continuity of their association. This approach effectively handles occlusions and temporary disappearances, ensuring robust tracking of shadow-object pairs throughout the video sequence.
IV Dataset and Evaluation Metrics
@@ IV-A Dataset
SOBA-VID comprises 292 videos totaling 7045 frames, sourced from (i) existing datasets, including DAVIS [44], YouTube-VIS [33], and OVIS [45], (ii) self-collected videos, and (iii) Internet searches using the keyword “shadow plus common instances.” We randomly partition the videos into a training set (232 videos, 5863 frames) and a testing set (60 videos, 1182 frames). For the test set videos, we meticulously annotate frame-by-frame masks of each shadow/object instance using Apple Pencil, forming a total of 134 shadow-object pairs (refer to Fig. 6 for examples). For the training set videos, we annotate one frame in every four frames, while ensuring that the last frame of each video is always annotated. This results in a total of 1719 annotated frames, containing 503 shadow-object pairs. Note that our semi-supervised framework does not use these annotations in Stage II, and we hope the labeled frames will promote future research in this field. Overall, our dataset contains 292 videos with a total of 7045 frames and 637 annotated shadow-object pairs. The average video length is 24.1 frames, and there are on average 2.2 shadow-object pairs per video.
IV-B Evaluation Metrics
@@
TABLE II: Comparison with three baseline methods for video instance shadow detection, showcasing our method’s substantial superiority. Improvements over “SSISv2 + Mask2Former” on SOAP-VID, Association AP, and Instance AP are 27.7%, 34.6%, and 31.2%, respectively.
@@
SSIS [2] + IoU Tracker[46] 21.5 31.7 25.7
SSIS [2] + Mask2Former[37] 29.2 42.5 36.2
SSISv2 [3] + Mask2Former[37] 31.0 45.7 38.8
@@
V Experimental Results
V-A Implementation Details
@@
For inference, ViShadow processes input videos online. Its tracking branch predicts tracking embeddings for both paired and unpaired shadow/object instances on a frame-by-frame basis, employing the matching strategy outlined in Section III-B. Notably, ViShadow demonstrates its efficiency by completing a 21-frame video in 18 seconds.
V-B Comparison with Baseline Methods
We compare our method with three baselines: (i) utilizing image-level detector SSIS [2] to detect the shadow instances, object instances, and shadow-object associations on each frame, and employing the IoU Tracker [46] to match instances across frames based on IoU values between detected bounding boxes; and (ii) employing the tracking algorithm Mask2former [37] to detect and track object instances across all video frames, using SSIS [2] for shadow/object instance and shadow-object association detection on each frame, merging the results based on mask IoU of each object instance, and similarly using SSISv2 [3] for the same task and merging the results in the same manner.
Table II [^1] The hyperparameters for the post-processing in SSIS + Mask2Former and SSISv2 + Mask2Former have been unified for comparison, leading to different results from the previous arXiv version. presents the comparison results, showing our method significantly outperforming both baselines across all evaluation metrics on the SOBA-VID test set. The IoU Tracker heavily relies on the quality of detection results and frame rate, while Mask2Former is constrained to recognizing objects of the same category as those in the training data [33]. In contrast, our approach benefits from end-to-end training and knowledge acquired from unlabeled videos. Visual comparisons between ViShadow and “SSIS + Mask2Former” are provided in Fig. 2.
@@ V-B Comparison with Baseline Methods
TABLE III: Evaluation on the instance paired tracking design.
@@
V-C Ablation Study
V-C1 Component Analysis
We conduct an ablation study, incrementally introducing major components—tracking branch with contrastive learning in Eq.(3), temporal correspondence in Eq.(4), and bidirectional retrieving mechanism—starting from the SSIS baseline, which employs instance tracking through the mask controller’s output. The results in Table I demonstrate the contributions of each component, with contrastive learning and temporal correspondence enhancing paired tracking, and the retrieving mechanism addressing missed unpaired instances.
V-C2 Instance Paired Tracking Analysis
We assess the effectiveness of instance paired tracking. Table III indicates that using tracking embeddings for individual shadow/object instances achieves paired tracking. And the best performance across all evaluation metrics is achieved by concatenating both shadow and object embeddings.
V-D Comparison with OmnimatteRF
@@
VI Applications
VI-A Mask-Guided Shadow-Object Manipulation
@@ VI-A3 Shadow Editing
VI-B Text-Instructed Shadow-Object Manipulation
@@
VII Discussion
VII-A Comparison to Other Video-Related Tasks
The tasks of video segmentation, video instance segmentation, and video shadow detection each address different aspects of video analysis but fall short in performing the specific tasks required for video instance shadow detection (VISD). Here’s an exploration of why these methods are inadequate for the VISD task and the unique requirements it entails.
Video segmentation focuses on partitioning a video into meaningful segments based on temporal and spatial features, often labeling different regions or objects within the frames. This task can be subdivided into semantic segmentation and instance segmentation. While semantic segmentation assigns a class label to each pixel, instance segmentation goes further by distinguishing between different instances of the same class. Note that the categories of pixels or instances are predefined. However, video segmentation methods do not identify or associate shadows with their corresponding objects. They treat shadows merely as background or part of the objects, which limits their ability to handle tasks that specifically require shadow-object pairing. Also, due to the limited categories, they cannot find all objects from the videos. Moreover, standard video segmentation methods do not track the temporal continuity of shadow-object pairs across frames, which is crucial for applications like shadow editing or video inpainting where maintaining the relationship between objects and their shadows over time is essential.
Video instance segmentation aims to detect, segment, and track individual object instances throughout video frames. This involves maintaining the identity of each object as it moves through the video sequence, leveraging both appearance and motion cues. While video instance segmentation methods excel at tracking objects, they only focus on objects with predefined categories, and they do not specifically address shadows. Additionally, these methods lack mechanisms to specifically detect and associate shadows with their respective objects, thus failing to support tasks that involve shadow-object interactions.
Video shadow detection focuses on identifying and segmenting shadows within video frames. It aims to distinguish shadows from other elements in the scene. However, video shadow detection methods do not identify each shadow instance and typically consider shadow regions as a single category. Consequently, they cannot establish the relationship between shadows and the objects that cast them. This limitation hinders their ability to handle tasks that require understanding the context of shadow-object pairs. Moreover, these methods often do not track shadows across frames, as they lack the concept of individual shadow instances.
Why these methods cannot perform the VISD task. VISD requires detecting, segmenting, associating, and tracking paired shadow-object instances in videos, which traditional methods cannot achieve. They lack the ability to establish clear shadow-object associations, maintain temporal continuity, and handle complex scenarios with occlusions and disappeared instances. Our framework, ViShadow, addresses these challenges with a semi-supervised approach using both labeled image data and unlabeled video data. It features a two-stage training pipeline and a retrieval mechanism for consistent tracking, enabling robust solutions for tasks like shadow removal, video inpainting, and shadow manipulation beyond the scope of traditional methods.
VII-B Limitations
This approach has certain limitations. First, distinguishing overlapping shadow instances from different objects is challenging because of their merged and indistinct boundaries. Second, accurately segmenting soft shadows, characterized by gradual edges and subtle intensity transitions, poses a significant challenge. Refer to Figure 11 for visual examples illustrating the challenges associated with overlapping and soft shadows. To address these limitations in future work, inspired by [59], we plan to use simulators, such as the game GTA-V, to render the overlapping and soft shadows, where the ground truth annotations can be generated automatically. This will expand the training and test datasets with more complex shadow interactions.
VIII Conclusion
@@
This paper introduces the video instance shadow detection task, aiming to detect, segment, associate, and track paired shadow-object instances in videos. We present the ViShadow framework, employing supervised learning on labeled images and self-supervised learning on unlabeled videos to learn paired tracking of shadow and object instances. A bidirectional retrieving mechanism is designed for robust paired tracking, even in cases where objects or shadows are temporarily missed. We contribute the SOBA-VID dataset for training and testing. Experimental results demonstrate the effectiveness of our approach across various metrics. Practical applications of video instance shadow detection in video-level editing scenarios are illustrated.
Acknowledgement
The work was supported by the National Key R&D Program of China (Grant No. 2023YFE0202700), the Research Grants Council of the Hong Kong Special Administrative Region, China (Grant No. 14201321), and the Hong Kong Innovation and Technology Fund (Grant No. MHP/092/22). Zhenghao Xing and Tianyu Wang are the joint first authors.